OpenStudy (anonymous):

(Equation Down Below in comments) 7) a. Additive Identity Property b. Additive Inverse Property Incorrect c. Multiplicative Identity Property d. Commutative Property of Multiplication e. Multiplicative Inverse Property 8) a. Symmetric Property of Equality b. Associative Property of Addition c. Distributive Property d. Commutative Property of Addition e. Associative Property of Multiplication

OpenStudy (anonymous):

7) \[\frac{ 2 }{ 3 } \times \frac{ 3 }{ 2 } = 1\]

OpenStudy (anonymous):

8) (a+4)+2=a+(4+2)